OpenStudy (anonymous):

Graph the equation. Describe the graph and its lines of symmetry. 4x^2 + 9y^2 = 36 The only thing I know about this is I think it's an ellipse. This is a problem from my conic sections class, please help!

OpenStudy (anonymous):

it is definitely an ellipse

OpenStudy (anonymous):

maybe it would be easier if you divided both sides by 36

OpenStudy (anonymous):

I've never tried that. The formula says that the 36 when square rooted is the radius, or so I thought, but an ellipse.

OpenStudy (anonymous):

\[\frac{x^2}{9}+\frac{y^2}{4}=1\] puts it in the form where you can read off what you need

OpenStudy (anonymous):

What is frac?

OpenStudy (anonymous):

this looks just like \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] where ok latex is not working for you refresh your browswer and you should see what i wrote

OpenStudy (anonymous):

I do not.

OpenStudy (anonymous):

the form you want it \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] so here you see that a =3 and b = 2

OpenStudy (anonymous):

can you read that?

OpenStudy (anonymous):

Yeah, thank you for your help!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you can also just cheat http://www.wolframalpha.com/input/?i=ellipse+4x^2+%2B+9y^2+%3D+36

OpenStudy (anonymous):

You're a lifesaver. This test is due today and I've got three problems I just can't understand!

OpenStudy (anonymous):

keep posting!

OpenStudy (anonymous):

I only have one more question, do you know of a website that is really good at showing you how to memorize equations, because often times after you learn to read an equation it become easier.